# Morita equivalence of Nijenhuis structures

### v1

@@
Abstract
We introduce Morita equivalence for Nijenhuis groupoids and for their infinitesimal counterparts, establishing a global-to-infinitesimal correspondence under the Lie functor. A special case is that of holomorphic Lie groupoids and algebroids. We use our framework to enhance the known Morita equivalences for quasi-symplectic groupoids and Dirac structures with compatible Nijenhuis structures.
1 Introduction
Nijenhuis structures provide a convenient setting for the study of integrability of several geometric structures. A key example is found in complex geometry, where a holomorphic structure on a manifold is equivalently described by an almost-complex structure satisfying the Nijenhuis condition (i.e., whose Nijenhuis torsion vanishes) [34, 18]. Nijenhuis structures are also important in the theory of integrable systems, serving as a tool for the systematic construction of hierarchies of functions in involution [29]. More recently, the study of Nijenhuis geometry as a subject in its own right has emerged (see, e.g. [4, 22, 5, 3]).
Nijenhuis structures on Lie groupoids naturally arise e.g. in the study of holomorphic Lie groupoids [25] and in the context of integration of Poisson-Nijenhuis manifolds to symplectic-Nijenhuis groupoids [36, 16, 15]. The primary goal of this paper is to establish an appropriate notion of Morita equivalence for Nijenhuis groupoids. This is achieved by introducing a general notion of Morita equivalence for multiplicative (1,1)-tensor fields which is compatible with the Nijenhuis condition, so it restricts to multiplicative Nijenhuis structures. This paves the way for the study of the Morita equivalence of more general quasi-Nijenhuis groupoids, a topic to be addressed in a future work.
@@ 1 Introduction
This framework finds a natural application in Poisson geometry. A compatibility between Poisson and Nijenhuis structures was given by F. Magri in [29] in the context of bihamiltonian integrable systems. This compatibility was later generalized to Dirac-Nijenhuis structures by Bursztyn-Drummond-Netto in [8] and integrated to its global counterpart, quasi-symplectic-Nijenhuis groupoids. We show that our definition of Morita equivalence of Nijenhuis structures preserves compatibility with these structures, at both the global and infinitesimal levels. This allows us to define a consistent notion of Morita equivalence for Dirac-Nijenhuis structures and quasi-symplectic-Nijenhuis groupoids. In particular, this includes the case of holomorphic presymplectic groupoids and holomorphic Dirac structures.
@@
  • Section 2: Following a review of Nijenhuis structures, we introduce Nijenhuis groupoids. We define Morita equivalence for multiplicative (1,1)-tensor fields (Definition 2.8). We prove in Theorem 2.11 that this constitutes an equivalence relation.
  • Section 3: We address the infinitesimal setting. We introduce Morita equivalence for infinitesimal Nijenhuis structures on Lie algebroids (Definition 3.8) and establish a correspondence with the global formulation. Theorem 3.18 and Theorem 3.19 establish the differentiation and integration procedures, respectively. This leads to Corollary 3.20, which defines a one-to-one correspondence between the global and infinitesimal Morita equivalence of Nijenhuis structures.
  • Section 4: We investigate compatibility conditions between the Morita equivalence of Nijenhuis structures and quasi-symplectic groupoids (globally) as well as twisted Dirac structures (infinitesimally). This yields an enhancement of the Morita equivalence of Dirac structures to Dirac-Nijenhuis structures. This encompasses, in particular, Morita equivalence of Poisson-Nijenhuis manifolds, as well that of as holomorphic Poisson and holomorphic Dirac structures. We establish the global-to-infinitesimal correspondence of these Morita equivalences: Theorem 4.17 addresses differentiation, and Theorem 4.18 covers integration.
@@ 2 Nijenhuis Groupoids
2.1 Preliminaries on Nijenhuis structures
We denote by $\Omega^{k}(P,TP)=\Gamma(\wedge^{k}T^{*}P\otimes TP)$ the space of $(1,k)$-tensor fields (also referred to as vector-valued $k$-forms) on a manifold $P$. Recall that given a $(1,1)$-tensor field $N\in\Omega^{1}(P,TP)$, its Nijenhuis torsion [35] is the $(1,2)$-tensor field $\mathscr{N}_{N}\in\Omega^{2}(P,TP)$ given by
$\mathscr{N}_{N}\left(u,v\right)\coloneqq\left[N(u),N(v)\right]-N\left(\left[N(u),v\right]+\left[u,N(v)\right]-N([u,v])\right),$ (1)
@@ 2.1 Preliminaries on Nijenhuis structures
Given a smooth map $\varphi:P\to B$, we say that two $(1,k)$-tensor fields $N\in\Omega^{k}(P,TP)$ and $N_{B}\in\Omega^{k}(B,TB)$ are $\varphi$-related (see e.g., [21]) if for every $p\in P$ and $u_{1},\ldots,u_{k}\in T_{p}P$ we have
@@
The Nijenhuis torsion is natural in the sense that $\varphi$-related (1,1)-tensor fields have $\varphi$-related Nijenhuis torsions; in particular given a $\varphi$-projectable $(1,1)$-tensor field $N\in\Omega^{1}(P,TP)$, its Nijenhuis torsion $\mathscr{N}_{N}\in\Omega^{2}(P,TP)$ is also $\varphi$-projectable [21, Section 8.15]. So if $N$ is a Nijenhuis $\varphi$-projectable $(1,1)$-tensor field $\varphi$-related to $N_{B}\in\Omega^{1}(B,TB)$, then the Nijenhuis torsion $\mathscr{N}_{N_{B}}$ of $N_{B}$ vanishes; hence $N_{B}$ is also a Nijenhuis tensor field.
@@
Consequently, the vector bundle $\varphi^{*}TB$ is isomorphic to $TP/\operatorname{ker}\mathrm{d}\varphi$. Recall that there is an action of the submersion groupoid $P\times_{B}P\rightrightarrows P$ on $\varphi^{*}TB$ given by $\left(p,q\right)\cdot\left(q,v\right)=\left(p,v\right)$, and it induces a unique action of $P\times_{B}P$ on $TP/\operatorname{ker}\mathrm{d}\varphi$ such that the isomorphism $\mathrm{d}\varphi$ is equivariant (see [27, Section 1.6]).
@@
Let $N\in\Omega^{1}(P,TP)$ be a $(1,1)$-tensor field and let $\varphi:P\to B$ be a surjective submersion. Then $N$ is $\varphi$-projectable if and only if $N\left(\operatorname{ker}\mathrm{d}\varphi\right)\subseteq\operatorname{ker}\mathrm{d}\varphi$ and the induced map $\bar{N}:TP/\operatorname{ker}\mathrm{d}\varphi\to TP/\operatorname{ker}\mathrm{d}\varphi$ is $(P\times_{B}P)$-equivariant.
@@
First, assume that $N$ is $\varphi$-projectable. There exists a tensor field $N_{B}\in\Omega^{1}(B,TB)$ such that $\mathrm{d}\varphi(N_{p}(u))=(N_{B})_{\varphi(p)}(\mathrm{d}\varphi(u))$ for every $u\in T_{p}P$. If $u\in\operatorname{ker}\mathrm{d}\varphi$, it follows that
@@
Thus, $N(\operatorname{ker}\mathrm{d}\varphi)\subseteq\operatorname{ker}\mathrm{d}\varphi$, and $N$ descends to a well-defined map $\bar{N}$ on the quotient bundle $TP/\operatorname{ker}\mathrm{d}\varphi$.
To establish equivariance, we use the fact that the isomorphism $\varphi^{*}TB\cong TP/\operatorname{ker}\mathrm{d}\varphi$ induces an action of $P\times_{B}P$ on the quotient bundle. Specifically, for an arrow $(p,q)\in P\times_{B}P$ acting on $\bar{v}\in(TP/\operatorname{ker}\mathrm{d}\varphi)_{q}$ to yield $\bar{u}\in(TP/\operatorname{ker}\mathrm{d}\varphi)_{p}$, the relation $(p,q)\cdot\bar{v}=\bar{u}$ holds if and only if $\mathrm{d}\varphi(u)=\mathrm{d}\varphi(v)$. Using the definition of the projected tensor $N_{B}$, we compute:
@@
Conversely, suppose that $N(\operatorname{ker}\mathrm{d}\varphi)\subseteq\operatorname{ker}\mathrm{d}\varphi$ and that the induced map $\bar{N}$ is equivariant. We define a tensor field $N_{B}\in\Omega^{1}(B,TB)$ by setting $(N_{B})_{b}(w)\coloneqq\mathrm{d}\varphi(N_{p}(u))$, where $p\in\varphi^{-1}(b)$ and $u\in T_{p}P$ satisfies $\mathrm{d}\varphi(u)=w$. The value of $(N_{B})_{b}(w)$ is independent of the choice of the lift $u$ because $N$ preserves the vertical kernel. Furthermore, it is independent of the choice of $p$ because the induced map $\bar{N}$ is $(P\times_{B}P)$-equivariant. Thus, $N_{B}$ is well-defined and $N$ is $\varphi$-projectable. ∎
@@
Corollary 2.2.
Suppose that a Lie groupoid $\mathscr{G}$ acts principally on $P$ with quotient map $\varphi:P\to B$. A $(1,1)$-tensor field $N\in\Omega^{1}(P,TP)$ is $\varphi$-projectable if and only if $N$ preserves the tangent spaces to the orbits of the action and the induced map $\bar{N}:TP/\operatorname{ker}\mathrm{d}\varphi\to TP/\operatorname{ker}\mathrm{d}\varphi$ is $\mathscr{G}$-equivariant.
@@
2.2 Multiplicative Nijenhuis structures
@@
is a Lie groupoid morphism. We denote by $\Omega^{k}(\mathscr{G},T\mathscr{G})$ the space of multiplicative $(1,k)$-tensor fields on $\mathscr{G}$. For details on multiplicative tensor fields we refer to [9].
@@
A Nijenhuis groupoid is a pair $(\mathscr{G},N)$, where $\mathscr{G}=\left(G_{1}\rightrightarrows G_{0}\right)$ is a Lie groupoid and $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ is a Nijenhuis tensor field; that is, $N$ is a multiplicative $(1,1)$-tensor field whose Nijenhuis torsion $\mathscr{N}_{N}$ vanishes.
@@
A holomorphic groupoid [25, 38] is a groupoid object in the category of complex manifolds; that is, the spaces of arrows and objects are complex manifolds and all structural maps are holomorphic. Equivalently, such a groupoid is a (real) Lie groupoid with an integrable multiplicative almost-complex structure [25, Proposition 3.16]. Thus, a holomorphic groupoid is a Nijenhuis groupoid $(\mathscr{G},I_{\mathscr{G}})$, where $I_{\mathscr{G}}\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ satisfies $I_{\mathscr{G}}^{2}=-\mathrm{id}_{T\mathscr{G}}$.
@@
Nijenhuis groupoids also arise naturally in the integration of Poisson-Nijenhuis manifolds (see Example 4.8). Stiénon and Xu [36] proved that if the underlying Poisson structure of a Poisson-Nijenhuis manifold $(M,\pi,r)$ integrates to a symplectic groupoid $(\mathscr{G},\omega)$, then $r$ lifts to a unique multiplicative Nijenhuis tensor field $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$. The compatibility between $\pi$ and $r$ ensures that $N$ and $\omega$ are compatible, endowing $\mathscr{G}$ with the structure of a symplectic-Nijenhuis groupoid.
The following lemma establishes the conditions under which a $(1,1)$-tensor field on a manifold $P$ is projectable when a Nijenhuis groupoid acts on $P$.
@@
Let $\mathscr{G}=(G_{1}\rightrightarrows G_{0})$ be a Lie groupoid and $\pi:P\to B$ a principal $\mathscr{G}$-bundle equipped with the action $\mathscr{A}:G_{1}\times_{G_{0}}P\to P$. Assume that $\mathscr{G}$ is equipped with a multiplicative $(1,1)$-tensor field $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ and that $P$ carries a $(1,1)$-tensor field $J\in\Omega^{1}(P,TP)$. If the action intertwines the tensor fields, i.e.,
$J\circ\mathrm{d}\mathscr{A}=\mathrm{d}\mathscr{A}\circ(N\times J),$ (2)
then $J$ projects to a unique $(1,1)$-tensor field $J_{B}\in\Omega^{1}(B,TB)$ on the base. Furthermore, if $N$ and $J$ are Nijenhuis tensor fields, then $J_{B}$ is also a Nijenhuis tensor field.
@@
We apply Corollary 2.2. First, we verify that $J$ preserves the vertical bundle $\operatorname{ker}\mathrm{d}\pi$, which coincides with the tangent spaces to the orbits of the $\mathscr{G}$-action. Let $\mathscr{O}$ be the orbit through a point $p\in P$, and write $x=\mu(p)$. For any $v\in T_{p}\mathscr{O}$, there exists $\xi\in\Gamma(A_{\mathscr{G}})$, where $A_{\mathscr{G}}$ is the Lie algebroid of $\mathscr{G}$, such that $v=\mathrm{d}\mathscr{A}_{(\mathbf{1}_{x},p)}\left(\overrightarrow{\xi}_{x},0_{p}\right)$. Applying condition (2) and the linearity of $J$ , we obtain:
@@
Since $N$ is a morphism of the tangent groupoid, it preserves the kernel of the source map differential. Consequently, there exists $\zeta\in\Gamma\left(A_{\mathscr{G}}\right)$ such that $N\left(\overrightarrow{\xi}\right)=\overrightarrow{\zeta}$. Therefore, $J(v)$ takes the form $J(v)=\mathrm{d}\mathscr{A}\left(\overrightarrow{\zeta}_{x},0_{p}\right)$, which implies $J(v)\in T_{p}\mathscr{O}$.
To establish that the induced map $\bar{J}\colon TP/\operatorname{ker}\mathrm{d}\pi\to TP/\operatorname{ker}\mathrm{d}\pi$ is $\mathscr{G}$-equivariant, we must show that $g\cdot\overline{J(v)}=\bar{J}(g\cdot\bar{v})$ for every $v\in T_{p}P$, and $(g,p)\in G_{1}\times_{G_{0}}P$. Under the identification $TP/\operatorname{ker}\mathrm{d}\pi\cong\varphi^{*}TB$, we may write $\bar{v}\cong(\varphi(p),\mathrm{d}\pi(v))$. Thus, $w\in T_{g\cdot p}P$ satisfies $g\cdot\bar{v}=\bar{w}$ if and only if $\mathrm{d}\pi(v)=\mathrm{d}\pi(w)$. We proceed as before: if $g\cdot\bar{v}=\bar{w}$, then there exists $\xi\in\Gamma(A_{\mathscr{G}})$ such that $w=\mathrm{d}\mathscr{A}\left(\overrightarrow{\xi}_{g},v\right)$. Applying condition (2), we obtain
@@
for some $\zeta\in\Gamma(A_{\mathscr{G}})$. Therefore, $\overline{J(w)}=g\cdot\overline{J(v)}$, which concludes the proof.
Finally, assume that $N$ and $J$ are Nijenhuis tensor fields. Since $J$ projects to $J_{B}$, the Nijenhuis torsion $N_{J}$ is $\pi$-related to $N_{J_{B}}$. The vanishing of $N_{J}$ implies the vanishing of $N_{J_{B}}$, proving that $J_{B}$ is a Nijenhuis tensor field. ∎
2.3 Morita equivalence of Nijenhuis groupoids
In order to introduce Morita equivalence of Nijenhuis groupoids, we recall two equivalent notions of Morita equivalence for Lie groupoids following [2, Theorem 2.2] (see also [32, Section 2.6]).

### v2

@@
Abstract.
We introduce Morita equivalence for Nijenhuis groupoids and for their infinitesimal counterparts, establishing a global-to-infinitesimal correspondence under the Lie functor. A special case is that of holomorphic Lie groupoids and algebroids. We use our framework to enhance the known Morita equivalences for quasi-symplectic groupoids and Dirac structures with compatible Nijenhuis structures. Finally, subject to certain conditions, we prove that the modular class of Poisson-Nijenhuis manifolds is invariant under Morita equivalence.
1. Introduction
Nijenhuis structures provide a convenient setting for the study of integrability of several geometric structures. A key example is found in complex geometry, where a holomorphic structure on a manifold is equivalently described by an almost-complex structure satisfying the Nijenhuis condition (i.e., whose Nijenhuis torsion vanishes) [undefah, undefr]. Nijenhuis structures are also important in the theory of integrable systems, serving as a tool for the systematic construction of hierarchies of functions in involution [undefab]. More recently, the study of Nijenhuis geometry as a subject in its own right has emerged (see, e.g. [undefd, undefv, undefe, undefc]).
Nijenhuis structures on Lie groupoids naturally arise e.g. in the study of holomorphic Lie groupoids [undefy] and in the context of integration of Poisson-Nijenhuis manifolds to symplectic-Nijenhuis groupoids [undefaj, undefp, undefo]. The primary goal of this paper is to establish an appropriate notion of Morita equivalence for Nijenhuis groupoids. This is achieved by introducing a general notion of Morita equivalence for multiplicative (1,1)-tensor fields which is compatible with the Nijenhuis condition, so it restricts to multiplicative Nijenhuis structures. This paves the way for the study of the Morita equivalence of more general quasi-Nijenhuis groupoids, a topic to be addressed in a future work.
@@ 1. Introduction
This framework finds a natural application in Poisson geometry. A compatibility between Poisson and Nijenhuis structures was established by Magri [undefab] in the context of bi-Hamiltonian integrable systems. This compatibility was later generalized to Dirac–Nijenhuis structures by Bursztyn, Drummond and Netto [undefh], and integrated to its global counterpart, quasi-symplectic–Nijenhuis groupoids. We show that our definition of Morita equivalence of Nijenhuis structures preserves compatibility with these structures, at both the global and infinitesimal levels. This allows us to define a consistent notion of Morita equivalence for Dirac–Nijenhuis structures and quasi-symplectic–Nijenhuis groupoids. In particular, this encompasses the case of holomorphic presymplectic groupoids and holomorphic Dirac structures.
We investigate the compatibility between Morita equivalence and the hierarchies of geometric structures naturally induced by Poisson–Nijenhuis and Dirac–Nijenhuis manifolds. Under suitable assumptions, we demonstrate that Morita equivalence persists throughout these hierarchies. As a cohomological consequence, we examine the intrinsic modular vector fields of Poisson–Nijenhuis manifolds introduced by Damianou and Fernandes [undefn]. Recall that the standard modular class of a Poisson manifold, which measures the obstruction to the existence of an invariant measure under Hamiltonian flows, is preserved under Morita equivalence [undefk]. In the enriched setting, the Poisson cohomology class of the intrinsic modular vector field serves as an additional modular class; its vanishing is the precise condition for the vector field to be bi-Hamiltonian. In Theorem 4.24, we establish that this modular class is also invariant under our introduced notion of Morita equivalence. Specifically, if two Poisson–Nijenhuis manifolds $(M_{1},\pi_{1},r_{1})$ and $(M_{2},\pi_{2},r_{2})$ are Morita equivalent via a non-degenerate infinitesimal bibundle, then their respective modular classes $\mathscr{Y}_{r_{1}}]$ and $\mathscr{Y}_{r_{2}}]$ correspond under the induced isomorphism between the first Poisson cohomology groups of $\big(M_{1},\pi_{1}^{(1)}\big)$ and $\big(M_{2},\pi_{2}^{(1)}\big)$.
@@
  • Section 2: Following a review of Nijenhuis structures, we introduce Nijenhuis groupoids and define Morita equivalence for multiplicative $(1,1)$-tensor fields (Definition 2.8). In Theorem 2.11, we prove that this notion constitutes an equivalence relation.
  • Section 3: Transitioning to the infinitesimal setting, we introduce Morita equivalence for infinitesimal Nijenhuis structures on Lie algebroids (Definition 3.8). Theorems 3.19 and 3.20 establish the respective differentiation and integration procedures, yielding a one-to-one correspondence between global and infinitesimal Morita equivalences (Corollary 3.21).
  • Section 4: We study the compatibility of the Morita equivalence for Nijenhuis structures with both quasi-symplectic groupoids and twisted Dirac structures. Consequently, we extend the established notions of Morita equivalence for quasi-symplectic groupoids and Dirac structures to quasi-symplectic-Nijenhuis groupoids and Dirac-Nijenhuis structures, respectively. We establish the global-to-infinitesimal correspondence via differentiation (Theorem 4.17) and integration (Theorem 4.18). Finally, in Theorem 4.24, we prove that the modular class of Poisson-Nijenhuis manifolds is invariant under Morita equivalence via non-degenerate infinitesimal bibundles.
@@ 2. Nijenhuis Groupoids
2.1. Preliminaries on Nijenhuis structures
We denote by $\Omega^{k}(P,TP)=\Gamma(\wedge^{k}T^{*}P\otimes TP)$ the space of $(1,k)$-tensor fields (also referred to as vector-valued $k$-forms) on a manifold $P$. Recall that given a $(1,1)$-tensor field $N\in\Omega^{1}(P,TP)$, its Nijenhuis torsion [undefai] is the $(1,2)$-tensor field $\mathscr{N}_{N}\in\Omega^{2}(P,TP)$ given by
(1) $\mathscr{N}_{N}\left(u,v\right)\coloneqq\left[N(u),N(v)\right]-N\left(\left[N(u),v\right]+\left[u,N(v)\right]-N([u,v])\right),$
@@ 2.1. Preliminaries on Nijenhuis structures
Given a smooth map $\varphi:P\to B$, we say that two $(1,k)$-tensor fields $N\in\Omega^{k}(P,TP)$ and $N_{B}\in\Omega^{k}(B,TB)$ are $\varphi$-related (see e.g., [undefu]) if for every $p\in P$ and $u_{1},\ldots,u_{k}\in T_{p}P$ we have
@@
The Nijenhuis torsion is natural in the sense that $\varphi$-related (1,1)-tensor fields have $\varphi$-related Nijenhuis torsions; in particular given a $\varphi$-projectable $(1,1)$-tensor field $N\in\Omega^{1}(P,TP)$, its Nijenhuis torsion $\mathscr{N}_{N}\in\Omega^{2}(P,TP)$ is also $\varphi$-projectable [undefu, Section 8.15]. So if NN is a Nijenhuis $\varphi$-projectable $(1,1)$-tensor field $\varphi$-related to $N_{B}\in\Omega^{1}(B,TB)$, then the Nijenhuis torsion $\mathscr{N}_{N_{B}}$ of $N_{B}$ vanishes; hence $N_{B}$ is also a Nijenhuis tensor field.
@@
Consequently, the vector bundle $\varphi^{*}TB$ is isomorphic to $TP/\operatorname{ker}\mathrm{d}\varphi$. Recall that there is an action of the submersion groupoid $P\times_{B}P\rightrightarrows P$ on $\varphi^{*}TB$ given by $\left(p,q\right)\cdot\left(q,v\right)=\left(p,v\right)$, and it induces a unique action of $P\times_{B}P$ on $TP/\operatorname{ker}\mathrm{d}\varphi$ such that the isomorphism $\mathrm{d}\varphi$ is equivariant (see [undefz, Section 1.6]).
@@
Let $N\in\Omega^{1}(P,TP)$ be a $(1,1)$-tensor field and let $\varphi:P\to B$ be a surjective submersion. Then NN is $\varphi$-projectable if and only if $N\left(\operatorname{ker}\mathrm{d}\varphi\right)\subseteq\operatorname{ker}\mathrm{d}\varphi$ and the induced map $\bar{N}:TP/\operatorname{ker}\mathrm{d}\varphi\to TP/\operatorname{ker}\mathrm{d}\varphi$ is $(P\times_{B}P)$-equivariant.
@@
First, assume that NN is $\varphi$-projectable. There exists a tensor field $N_{B}\in\Omega^{1}(B,TB)$ such that $\mathrm{d}\varphi(N_{p}(u))=(N_{B})_{\varphi(p)}(\mathrm{d}\varphi(u))$ for every $u\in T_{p}P$. If $u\in\operatorname{ker}\mathrm{d}\varphi$, it follows that
@@
Thus, $N(\operatorname{ker}\mathrm{d}\varphi)\subseteq\operatorname{ker}\mathrm{d}\varphi$, and NN descends to a well-defined map $\bar{N}$ on the quotient bundle $TP/\operatorname{ker}\mathrm{d}\varphi$.
To establish equivariance, we use the fact that the isomorphism $\varphi^{*}TB\cong TP/\operatorname{ker}\mathrm{d}\varphi$ induces an action of $P\times_{B}P$ on the quotient bundle. Specifically, an arrow $(p,q)\in P\times_{B}P$ acts on $\bar{v}\in(TP/\operatorname{ker}\mathrm{d}\varphi)_{q}$ to yield $(p,q)\cdot\bar{v}=\bar{u}\in(TP/\operatorname{ker}\mathrm{d}\varphi)_{p}$ if and only if $\mathrm{d}\varphi(u)=\mathrm{d}\varphi(v)$. Using the definition of the projected tensor $N_{B}$, we compute:
@@
Conversely, suppose that $N(\operatorname{ker}\mathrm{d}\varphi)\subseteq\operatorname{ker}\mathrm{d}\varphi$ and that the induced map $\bar{N}$ is equivariant. We define a tensor field $N_{B}\in\Omega^{1}(B,TB)$ by setting $(N_{B})_{b}(w)\coloneqq\mathrm{d}\varphi(N_{p}(u))$, where $p\in\varphi^{-1}(b)$ and $u\in T_{p}P$ satisfies $\mathrm{d}\varphi(u)=w$. The value of $(N_{B})_{b}(w)$ is independent of the choice of the lift $u$ because NN preserves the vertical kernel. Furthermore, it is independent of the choice of $p$ because the induced map $\bar{N}$ is $(P\times_{B}P)$-equivariant. Thus, $N_{B}$ is well-defined and NN is $\varphi$-projectable. ∎
@@
Corollary $2.$
Suppose that a Lie groupoid $\mathscr{G}$ acts principally on $P$ with quotient map $\varphi:P\to B$. A $(1,1)$-tensor field $N\in\Omega^{1}(P,TP)$ is $\varphi$-projectable if and only if NN preserves the tangent spaces to the orbits of the action and the induced map $\bar{N}:TP/\operatorname{ker}\mathrm{d}\varphi\to TP/\operatorname{ker}\mathrm{d}\varphi$ is $\mathscr{G}$-equivariant.
@@
$2.$ Multiplicative Nijenhuis structures
@@
is a Lie groupoid morphism. We denote by $\Omega^{k}(\mathscr{G},T\mathscr{G})$ the space of multiplicative $(1,k)$-tensor fields on $\mathscr{G}$. For details on multiplicative tensor fields we refer to [undefg].
@@
A Nijenhuis groupoid is a pair $(\mathscr{G},N)$, where $\mathscr{G}=\left(G_{1}\rightrightarrows G_{0}\right)$ is a Lie groupoid and $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ is a Nijenhuis tensor field; that is, NN is a multiplicative $(1,1)$-tensor field whose Nijenhuis torsion $\mathscr{N}_{N}$ vanishes.
@@
A holomorphic groupoid [undefy, undefal] is a groupoid object in the category of complex manifolds; that is, the spaces of arrows and objects are complex manifolds and all structural maps are holomorphic. Equivalently, such a groupoid is a (real) Lie groupoid with an integrable multiplicative almost-complex structure [undefy, Proposition 3.16]. Thus, a holomorphic groupoid is a Nijenhuis groupoid $(\mathscr{G},I_{\mathscr{G}})$, where $I_{\mathscr{G}}\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ satisfies $I_{\mathscr{G}}^{2}=-\mathrm{id}_{T\mathscr{G}}$.
@@
Nijenhuis groupoids also arise naturally in the integration of Poisson-Nijenhuis manifolds (see Example 4.8). Stiénon and Xu [undefaj] proved that if the underlying Poisson structure of a Poisson-Nijenhuis manifold $(M,\pi,r)$ integrates to a symplectic groupoid $(\mathscr{G},\omega)$, then $r$ lifts to a unique multiplicative Nijenhuis tensor field $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$. The compatibility between $\pi$ and $r$ ensures that NN and $\omega$ are compatible, endowing $\mathscr{G}$ with the structure of a symplectic-Nijenhuis groupoid.
The following lemma establishes conditions under which a $(1,1)$-tensor field on a manifold $P$ is projectable when a Nijenhuis groupoid acts on $P$.
@@
Let $\mathscr{G}=(G_{1}\rightrightarrows G_{0})$ be a Lie groupoid and $\pi:P\to B$ a principal $\mathscr{G}$-bundle equipped with an action $\mathscr{A}:G_{1}\times_{G_{0}}P\to P$. Assume that $\mathscr{G}$ is equipped with a multiplicative $(1,1)$-tensor field $N\in\Omega^{1}(\mathscr{G},T\mathscr{G})$ and that $P$ carries a $(1,1)$-tensor field $J\in\Omega^{1}(P,TP)$. If the action intertwines the tensor fields, i.e.,
(2) $J\circ\mathrm{d}\mathscr{A}=\mathrm{d}\mathscr{A}\circ(N\times J),$
then $J$ projects to a unique $(1,1)$-tensor field $J_{B}\in\Omega^{1}(B,TB)$ on the base. Furthermore, if NN and $J$ are Nijenhuis tensor fields, then $J_{B}$ is also a Nijenhuis tensor field.
@@
We apply Corollary $2.$ First, we verify that $J$ preserves the vertical bundle $\operatorname{ker}\mathrm{d}\pi$, which coincides with the tangent spaces to the orbits of the $\mathscr{G}$-action. Let $\mathscr{O}$ be the orbit through a point $p\in P$, and write $x=\mu(p)$. For any $v\in T_{p}\mathscr{O}$, there exists $\xi\in\Gamma(A_{\mathscr{G}})$, where $A_{\mathscr{G}}$ is the Lie algebroid of $\mathscr{G}$, such that $v=\mathrm{d}\mathscr{A}_{(\mathbf{1}_{x},p)}\left(\overrightarrow{\xi}_{x},0_{p}\right)$. Applying condition (2) and the linearity of $J$, we obtain:
@@
Since NN is a morphism of the tangent groupoid, it preserves the kernel of the source map differential. Consequently, there exists $\zeta\in\Gamma\left(A_{\mathscr{G}}\right)$ such that $N\left(\overrightarrow{\xi}\right)=\overrightarrow{\zeta}$. Therefore, $J(v)$ takes the form $J(v)=\mathrm{d}\mathscr{A}\left(\overrightarrow{\zeta}_{x},0_{p}\right)$, which implies $J(v)\in T_{p}\mathscr{O}$.
To establish that the induced map $\bar{J}\colon TP/\operatorname{ker}\mathrm{d}\pi\to TP/\operatorname{ker}\mathrm{d}\pi$ is $\mathscr{G}$-equivariant, we must show that $g\cdot\bar{J}(\bar{v})=\bar{J}(g\cdot\bar{v})$ for every $v\in T_{p}P$, and $(g,p)\in G_{1}\times_{G_{0}}P$. Under the identification $TP/\operatorname{ker}\mathrm{d}\pi\cong\varphi^{*}TB$, we may write $\bar{v}\cong(\varphi(p),\mathrm{d}\pi(v))$. Thus, $w\in T_{g\cdot p}P$ satisfies $g\cdot\bar{v}=\bar{w}$ if and only if $\mathrm{d}\pi(v)=\mathrm{d}\pi(w)$. We proceed as before: if $g\cdot\bar{v}=\bar{w}$, then there exists $\xi\in\Gamma(A_{\mathscr{G}})$ such that $w=\mathrm{d}\mathscr{A}\left(\overrightarrow{\xi}_{g},v\right)$. Applying condition (2), we obtain
@@
for some $\zeta\in\Gamma(A_{\mathscr{G}})$. Therefore, $\bar{J}(\bar{w})=g\cdot\overline{J(v)}$, which concludes the proof.
Finally, assume that NN and $J$ are Nijenhuis tensor fields. Since $J$ projects to $J_{B}$, the Nijenhuis torsion $N_{J}$ is $\pi$-related to $N_{J_{B}}$. The vanishing of $N_{J}$ implies the vanishing of $N_{J_{B}}$, proving that $J_{B}$ is a Nijenhuis tensor field. ∎
2.3. Morita equivalence of Nijenhuis groupoids
In order to introduce Morita equivalence of Nijenhuis groupoids, we recall two equivalent notions of Morita equivalence for Lie groupoids following [undefb, Theorem 2.2] (see also [undefae, Section 2.6]).